OpenStudy (anonymous):

Seetharamu has a mass of 60kg. He is wearing a pair of roller skates at rest and holding a bog of mass 5kg.When he throws the back fowards at a velocity if 4ms-1,he moves at a velocity of vms-1. Determine the a) respective initial momentum of Seetharamu and the bag, b)momentum of the bag right after being thrown, c)value of v.

OpenStudy (ajprincess):

Initial momentum of Seetharamu and the bag is zero because it is given that he is at rest initially. is that k?

OpenStudy (anonymous):

Yes.

OpenStudy (ajprincess):

Nw the bag is thrown with the velocity 4ms^-1. as u knw its momentum will be its mass*velocity,

OpenStudy (anonymous):

ok.

OpenStudy (ajprincess):

Then p=5*4 =20kgms^-1

OpenStudy (anonymous):

How about the 'c' ? I dont understand that actually.

OpenStudy (ajprincess):

Nw let's calculate the momentum of Seetharamu. It will b his mass multiplied by the velocity. That is p1=60v.

OpenStudy (ajprincess):

do u knw abt conservation of momentum?

OpenStudy (anonymous):

no :(

OpenStudy (ajprincess):

total linear momentum of system of body, if no external forces are acting, remains constant. That is, Total initial momentum = total final momentum

OpenStudy (ajprincess):

is that k?

OpenStudy (anonymous):

ok

OpenStudy (ajprincess):

Nw final momentum of Seetharamu and the bag = 20+60v Initial momentum = 0 nw equating both of them gives u 0=20+60v 60v=-20 v=-20/60 v=-1/3. This is in minus because I assumed that both of them are moving in the same direction.

OpenStudy (ajprincess):

Suppose there are two masses m1 and m2 moving in the same direction vth the velocities v1 and v2 respectively. aftr a certain tym they collide and move vth the velocities v3 and v4 respectively.

OpenStudy (ajprincess):